# zkMixer: A Configurable Zero-Knowledge Mixer with Proof of Innocence and Anti-Money Laundering Consensus Protocols

### v1

@@ Abstract
mixer, zero-knowledge, proof of innocence, cryptocurrencies, blockchain, privacy
1 Introduction
@@
Most cryptocurrencies, such as Bitcoin, were created to facilitate the exchange of value, i.e., act as money. However, the transparent and immutable nature of blockchains means that most blockchain systems only provide pseudo-anonymity, which offers less privacy than traditional cash.
@@
There are many ways for users to increase their privacy when transacting using cryptocurrencies, each with varying beneficial and detrimental properties. Common approaches include using privacy-focused cryptocurrencies, such as Monero [1] and Zcash [2]; moving funds through different wallets (wallet hopping) and decentralized finance (DeFi) applications to complicate transaction tracing; moving funds between different blockchains using bridges; using centralized services such as over-the-counter (OTC) brokers or exchanges [3]; and interacting with mixing services.
@@
Mixing services – also known as mixers, tumblers, or blenders – are common in the blockchain space and are designed to increase privacy by breaking the linkability of transactions. As the name suggests, mixers offer privacy by mixing the funds of many users, thus granting additional anonymity by adding users to an anonymity group of all the other users of the mixing service. Unfortunately, the anonymity provided by mixers means that they are commonly used by criminals looking to launder illicit funds. Since 2017, North Korean hackers, alone, have stolen more than \$6B of crypto assets, with much of this laundered using mixers and other layering techniques [4]. [^1] While we were writing this paper (21 Feb 2025), North Korean hackers conducted the largest known theft of any kind, stealing \$1.46B from Bydit, a Dubai-based centralized cryptocurrency exchange [4].
@@
In this paper, we tackle the problem of money laundering through mixers by introducing a novel mixer protocol that uses zero-knowledge proofs (ZKP) and a consensus process to monitor mixer input and decide whether to allow, refuse, or confiscate deposits. Our protocol enables users to establish complex criteria on who is allowed to deposit funds to the mixer, while preserving user/deposit anonymity.
2 Background
2.1 Illicit Activity on the Blockchain
@@
The blockchain analysis company, Chainalysis, has concretely identified \$40.9 billion of illicit on-chain activity in 2024, and estimates that the true amount could be closer to \$51 billion [5], which corresponds to 0.14% of the total annual transaction volume. Most of the illicit activity identified (63%) involved the use of stablecoins, and included transfers to sanctioned entities/jurisdictions (\$15.8B), cybercrime (\$10.8B), frauds and scams (\$9.9B), market manipulation (\$2.6B), stolen funds (\$2.2B), darknet markets (\$2B), and ransomware (\$0.8B).
@@
Mixers are not immune from crime, and their privacy-enhancing properties can attract bad actors who are looking to launder stolen funds or proceeds of crime. Chainalysis tracked \$31.5B (2022) and \$22.2B (2023) of laundered cryptocurrencies [6]. Of these, \$1.01B (2022) and \$504.3M (2023) were moved through mixers. To put this into perspective, in 2022 the total amount moved through mixers was \$7.8B [7], and Tornado Cash, which used to be one of the most popular mixers, had more than 25% of its deposits linked to illicit activities [5].
@@
Mixers therefore make cryptocurrency crime easier and more prevalent. Criminals using mixers also degrade the experience of mixers for legitimate users. Legitimate users who only wish to use mixers to gain privacy (e.g., to secretly donate to a cause that is frowned upon in their country [8]), must accept the stigma of holding funds that are potentially associated with criminal activities. Even worse, some individuals, businesses, or even smart contracts may choose not to accept funds originating from certain mixers [9]. This limits the use of mixer-anonymized cryptocurrencies, and in some sense, funds that are passed through a mixer become tainted.
@@
To combat crime, the governments of the US and several European countries have started to take down servers that host mixers, arrest their operators, and even sanction the mixers themselves. Notable server seizures include Bestmixer.io [10], ChipMixer [11], and Sinbad.io [12]. Notable arrests include the operators of Bitcoin Fog [13], Helix [14], Samourai Wallet [15] and the creators of Tornado Cash [16, 17]. The arrest of the founders of Tornado Cash is interesting because, unlike the other examples presented here, Tornado Cash is a decentralized smart contract that the founders do not directly control. The Office of Foreign Assets Control (OFAC) of the US Treasury has also sanctioned several mixers, including Blender.io [18], Tornado Cash [19, 20], and Sinbad.io [21]. Again, the case of Tornado Cash is interesting, as sanctions were later overturned by the US Court of Appeals, stating, “we hold that Tornado Cash’s immutable smart contracts (the lines of privacy-enabling software code) are not the ‘property’ of a foreign national or entity, meaning (1) they cannot be blocked under IEEPA, and (2) OFAC overstepped its congressionally defined authority” [22]. Funds originating from a sanctioned mixer can also be used for a malicious purpose, such that individuals are sent small amounts of these funds – a process known as “dusting” – in order to put the receivers of funds in breach of sanctions.
2.2 Types of Mixers
@@
Decentralized mixers do not depend on an operator but on cryptography. Specifically, most of these mixers either use transactions with multiple inputs and outputs – an idea first introduced in early 2013 by Maxwell and known as CoinJoins [23, 24] – or Zero-knowledge proofs (ZKPs) [25]. A CoinJoin on its own requires a coordinator, but this can be avoided by using a peer-to-peer network. However, in both cases, the coordinator or the rest of the peers have full knowledge of all transaction routing. This can be prevented by using blind signatures [26] or more complex protocols such as CoinShuffle [27] and CoinShuffle++ [28]. CoinJoins are simpler than ZKPs and are mainly used in blockchains such as Bitcoin, which do not support complex operations. However, the anonymity set is usually limited. Zero-knowledge (ZK) based mixers provide better anonymity, as their anonymity sets include all users of the mixer, and not just those of a single CoinJoin instance. However, ZKPs require more complex operations and are therefore only suitable for smart contract-enabled blockchains such as Ethereum.
@@
2.3 Zero-Knowledge Proofs
A zero-knowledge proof (ZKP) allows a party (the prover) to prove a statement to another party (the verifier) without revealing anything apart from the statement that is being proven. Although early ZKPs were impractical, recent improvements, such as the introduction of zero-knowledge succinct non-interactive arguments of knowledge (zk-SNARKs) [29], have enabled practical applications (see Nitulescu [30] for a history of SKP systems). Today, zk-SNARK variations are commonly used in decentralized blockchain-based systems, e.g., Pinocchio [31], Groth16 [32], Bulletproofs [33], and PlonK [34] (see Liang et al. [35] for a survey of common systems and tools for zk-SNARKs).
2.4 Zero-Knowledge Cryptocurrency Mixers
ZK cryptocurrency mixers work by breaking the link between the sender and the recipient of a transaction. The main idea is that, instead of sending funds directly to a recipient, the sender first sends the funds to a smart contract and allows the recipient to later withdraw them. Of course, implementing this naively, e.g., by using a mapping holding user balances, would not break the link between the sender and the recipient; it would at best make observing this connection slightly more difficult. This is where ZKPs enter the picture. Using ZKPs, a system can be designed so that there is no way to link the input with the output of a transaction passing through the mixer smart contract. To explain how this works, we will describe a simplified version of Tornado Cash [36], as most ZK cryptocurrency mixers follow a similar design (see Fig. 2).
@@ 2.4 Zero-Knowledge Cryptocurrency Mixers
Additional tools and processes can be used to further improve the user experience or increase compliance and anonymity. Examples are transaction relayers [38] and viewing keys [39]. In this paper, we will not discuss any of these tools as they are not integral to our protocol and are not used by our prototype. However, there is no reason why our protocol cannot be made compatible with such tools.
2.5 Proof of Innocence
@@
In 2023, a new method, called Proof of Innocence (PoI), was introduced to act as a compliance system for mixers [40, 41]. PoI is an extra layer on top of a mixer that allows a user to specify a list of existing deposits and prove that their withdrawal does not correspond to any of them. Privacy Pools [42], is a similar protocol that allows both inclusion and exclusion proofs. If this list contains all deposits to the mixer that are related to illegal activities, PoI should allow a user to prove that their withdrawal is not linked to proceeds of crime. In contrast, when a malicious user tries to launder funds that come from illicit activities, they will be unable to generate a valid PoI. We show an example of PoI failure in Fig. 3, scenario 1.
However, PoI does not provide the expected security guarantees, as it is simple to bypass the PoI system by using the mixer itself to first “clean” funds before requesting PoI validation. We show an example of PoI bypass in Fig. 3, scenario 2. The process is described in the following:
@@
    At this point, $M_{2}$ would not be able to use PoI, as $C_{1}$ is in $L$ (i.e., $M_{2}$’s withdrawal would correspond to a deposit that is linked to an illegal activity)
@@
    $M_{2}$’s withdrawal is indistinguishable from any other output from the mixer (as long as PoI is not applied) and so cannot be classified as malicious
@@
    $M_{2}$ then sends their new funds to the mixer and creates a new deposit with commitment $C_{2}$
@@
    Finally, $M_{3}$ generates a valid PoI, as their withdrawal with commitment $C_{2}$ does not correspond to any of the deposits contained in list $L$
@@ 2.6 Research Aim
We have identified that mixers are commonly used by criminals to launder proceeds of crime (Section 2.1); and we have shown that current PoI systems are insufficient, as they can be easily defeated (Section 2.5). Therefore, if we want to keep the benefits of anonymity provided by cryptocurrency mixers but also avoid their abuse by bad actors, it is clear that a new design is necessary. To achieve this, we propose a novel design that integrates innocence checks into the mixer (rather than as a separate component, as with current PoI systems); all checks are performed before funds are added to the mixer; and all deposits are verified. This approach ensures that all mixer outputs can be trusted; eliminating the threat of attackers using the mixer to trick the PoI system (as shown in Fig. 3).
3 System Design and Implementation
3.1 Improvements on ZK Mixer Design
@@
Our protocol improves the design of mixers in three main ways. First, we prevalidate the deposits before they enter the mixer to ensure that they adhere to a set of predefined rules that are configured when the mixer instance is first created. Rules can range from very simple checks that can be performed automatically by smart contracts, such as “only admit inputs from users who hold a soulbound token (SBT)” [46], to more complex schemes that check the origin of funds or check a user’s historical interactions with other mixers. These latter rules are too complex for a smart contract to perform, and so a group of validators is required. The validators check every input to the mixer and ensure that it follows the rules of the specific protocol instance. This makes the system highly customizable and eliminates the PoI issue (described earlier in Section 2.5).
@@
Second, we introduce a timer delay lock before tokens are added to the Merkel tree of the mixer. The delay is designed to provide the validators with time to verify each input: it can be set as a constant, e.g., one day; or it can be set to vary dynamically, e.g., increasing with system traffic (see Fig. 4 for available delay modes). To be effective, a constant delay must be long enough to ensure that malicious activity can be detected during periods of high traffic; while a variable delay enables both efficient mixer operation during low traffic and sufficient time for validation during high traffic. High traffic bursts over a short period can themselves be an indication of laundering (see, e.g., the heuristics identified in [47]), so dynamically slowing mixer operation during these times acts as a further safety valve, as an unusually large deposit will trigger a marked increase in the delay time, which flags anomalous deposit behaviour.
@@
Third, we allow validators to control inputs to the mixer during the lock period. Using a common consensus protocol, validators have the option of admitting deposits, rejecting and returning deposits to the sender, or confiscating deposits. Although confiscation should be used to restitute funds to victims of crime, it can be abused if the validators collude. Therefore, when a system instance is first deployed, the confiscation functionality can be enabled or disabled: when disabled, the validators only have the option to admit the deposit or reject and return the deposit to the sender. Here, we must stress again that validators can only assume control over deposits during the lock period. Once a deposit is accepted into the mixer, it is controlled by the user as in any other ZK-based mixer. Many smart contracts commonly used in DeFi have similar, or even stronger, control mechanisms. For example, the issuers of the most popular stablecoins such as USDC, USDT, and BUSD have the power to freeze or burn stablecoins and do so regularly when ordered by authorities [48]. This property of our system might not satisfy maximalist notions of decentralization, but we believe that it is sufficient for practical use in most cases.
3.2 System Components

### v2

# zkMixer: A Configurable Zero-Knowledge Mixer with Anti-Money Laundering Consensus Protocols

@@ Abstract
cryptocurrency mixer, zero-knowledge, proof-of-innocence, cryptocurrencies, blockchain technology, privacy
I Introduction
@@
Most cryptocurrencies, such as Bitcoin, were created to facilitate the exchange of value, i.e., act as money. However, the transparent and immutable nature of blockchains means that most blockchain systems only provide pseudo-anonymity, which is less anonymity than traditional cash.
@@
There are many ways for users to increase their privacy when transacting using cryptocurrencies, each with varying beneficial and detrimental properties. Common approaches include using privacy-focused cryptocurrencies, such as Monero [1] and Zcash [2]; moving funds through different wallets (wallet hopping) and decentralized finance (DeFi) applications to complicate transaction tracing; moving funds between different blockchains using bridges; using centralized services such as over-the-counter (OTC) brokers or exchanges; and interacting with mixing services.
@@
Mixing services – also known as mixers, tumblers, or blenders – are common in the blockchain space and are designed to increase privacy by breaking the linkability of transactions. As the name suggests, mixers offer anonymity by mixing the funds of many users. This grants anonymity to a user by adding them to the anonymity set of all the users of the mixing service. Unfortunately, the anonymity provided by mixers means that they are commonly used by criminals looking to launder illicit funds. Since 2017, North Korean hackers have stolen more than \$6 billion of crypto assets, with much of this laundered using mixers and other layering techniques [3]. [^1] While we were writing this paper (21 Feb 2025), North Korean hackers conducted the largest known theft of any kind, stealing \$1.46 billion from Bybit, a Dubai-based centralized cryptocurrency exchange [3].
@@
In this paper, we tackle the problem of money laundering through mixers by introducing a novel mixer protocol that uses zero-knowledge proofs (ZKP) and a consensus process to monitor mixer inputs and decide whether to allow, refuse, or confiscate deposits. Our protocol enables users to establish complex criteria on who is allowed to deposit funds to the mixer, while preserving user/deposit anonymity.
II Background
II-A Illicit Activity on the Blockchain
@@
The blockchain analysis company, Chainalysis, has concretely identified \$40.9 billion of illicit on-chain activity in 2024, and estimates that the true amount could be closer to \$51 billion [4], which corresponds to 0.14% of the total annual transaction volume. Most of the illicit activity identified (63%) involved the use of stablecoins, and included transfers to sanctioned entities/jurisdictions (\$15.8B), cybercrime (\$10.8B), frauds and scams (\$9.9B), market manipulation (\$2.6B), stolen funds (\$2.2B), darknet markets (\$2B), and ransomware (\$0.8B) [4].
@@
Mixers are not immune from crime, and their privacy-enhancing properties can attract bad actors who are looking to launder stolen funds or proceeds of crime. Chainalysis tracked \$31.5B (2022) and \$22.2B (2023) of laundered cryptocurrencies [5]. Of these, \$1.01B (2022) and \$0.5B (2023) were moved through mixers. To put this into perspective, in 2022 the total amount moved through mixers was \$7.8B [6], and Tornado Cash, which used to be one of the most popular mixers, had more than 25% of its deposits linked to illicit activities [4].
@@
Criminals using mixers also degrade the experience of mixers for legitimate users. Legitimate users, who only wish to use mixers to gain anonymity, must accept the stigma of holding funds that are potentially associated with criminal activities. Even worse, some individuals, businesses, or even smart contracts may choose not to accept funds originating from certain mixers [7]. This limits the use of mixer-anonymized cryptocurrencies and, in some sense, funds that are passed through a mixer become tainted.
@@
To combat crime, the governments of the US and several European countries have started to take down servers that host mixers, arrest their operators, and even sanction the mixers themselves. Notable server seizures include Bestmixer.io [8], ChipMixer [9], and Sinbad.io [10]. Notable arrests include the operators of Bitcoin Fog [11], Helix [12], Samourai Wallet [13], and the creators of Tornado Cash [14]. The arrest of the founders of Tornado Cash is interesting because, unlike the other examples presented here, Tornado Cash is a decentralized smart contract that the founders do not directly control. The Office of Foreign Assets Control (OFAC) of the US Treasury has also sanctioned several mixers, including Blender.io [15], Tornado Cash [16], and Sinbad.io [17]. Again, the case of Tornado Cash is interesting, as sanctions were later overturned by the US Court of Appeals, stating, “we hold that Tornado Cash’s immutable smart contracts (the lines of privacy-enabling software code) are not the ‘property’ of a foreign national or entity, meaning (1) they cannot be blocked under IEEPA, and (2) OFAC overstepped its congressionally defined authority” [18]. Funds originating from a sanctioned mixer can also be used for a malicious purpose, such that individuals are sent small amounts of these funds – a process known as “dusting” – in order to put the receivers of funds in breach of sanctions.
II-B Types of Mixers
@@
Decentralized mixers do not depend on an operator but on cryptography. Specifically, most of these mixers either use transactions with multiple inputs and outputs — an idea first introduced in early 2013 by Maxwell and known as CoinJoins [19, 20] — or Zero-knowledge proofs (ZKPs) [21]. A CoinJoin on its own requires a coordinator, but this can be avoided by using a peer-to-peer network. However, in both cases, the coordinator, or the rest of the peers, have complete knowledge of all transaction routing. This can be prevented by using blind signatures [22] or more complex protocols, such as CoinShuffle [23]. CoinJoins are simpler than ZKPs and are mainly used in blockchains such as Bitcoin, which do not support complex operations. However, the anonymity set is usually limited. Zero-knowledge (ZK) based mixers provide better anonymity, as their anonymity sets include all users of the mixer, and not just those of a single CoinJoin instance. However, ZKPs require more complex operations and are therefore only suitable for smart contract-enabled blockchains such as Ethereum.
@@
II-C Zero-Knowledge Proofs
A zero-knowledge proof (ZKP) allows a party (the prover) to prove a statement to another party (the verifier) without revealing anything apart from the statement that is being proven. Although early ZKPs were impractical for use in smart contracts, recent improvements, such as the introduction of zero-knowledge succinct non-interactive arguments of knowledge (zk-SNARKs) [24], have enabled practical blockchain applications (see Nitulescu [25] for a history of ZKP systems). Today, zk-SNARK variations are commonly used in decentralized blockchain-based systems, e.g., Pinocchio [26], Groth16 [27], and PlonK [28] (see Liang et al. [29] for a survey of common systems and tools for zk-SNARKs).
II-D Zero-Knowledge Cryptocurrency Mixers
ZK cryptocurrency mixers work by breaking the link between the sender and the recipient of a transaction. The main idea is that, instead of sending funds directly to a recipient, the sender first sends the funds to a smart contract and allows the recipient (which could be a fresh address controlled by the sender) to later withdraw them. Of course, implementing this naively, e.g., by using a mapping holding user balances, would not break the link between the sender and the recipient; it would at best make observing this connection slightly more difficult. This is where ZKPs enter the picture. Using ZKPs, a system can be designed so that there is no way to link the input with the output of a transaction passing through the mixer smart contract. To explain how this works, we will describe a simplified version of Tornado Cash [30], as most ZK cryptocurrency mixers follow a similar design (see Fig. 2).
@@ II-D Zero-Knowledge Cryptocurrency Mixers
Additional tools and processes can be used to further improve the user experience or increase compliance and anonymity. Examples are transaction relayers [32] and viewing keys [33]. In this paper, we will not discuss any of these tools, as they are not integral to our protocol and are not used by our prototype. However, there is no reason why our protocol cannot be made compatible with such tools.
II-E Proof of Innocence
@@
In late 2022, a new method called Proof of Innocence (PoI) was introduced to act as a compliance system for mixers [34, 35]. PoI is an extra voluntary layer on top of a mixer that allows a user to specify a list of existing deposits and then prove that their withdrawal does not correspond to any of them. In practice, these lists are usually compiled by blockchain analytics companies that provide real-time screening services. [^2] For example, Chainlink provides such a service: https://www.chainalysis.com/free-cryptocurrency-sanctions-screening-tools/ If this list contains all deposits to the mixer that are related to illegal activities, PoI should allow a user to prove that their withdrawal is not linked to proceeds of crime. In contrast, when a malicious user tries to launder funds that come from illicit activities, they will be unable to generate a valid PoI. In Fig. 3, Scenario 1, we show an example of a correct PoI failure, because the deposit with commitment $C_{1}$ is included in the list of illicit funds, $L$.
II-E1 Attack on PoI
PoI does not provide the expected security guarantees, as it is simple to bypass the PoI system by using the mixer itself to first “clean” funds before requesting PoI validation. We show an example of PoI bypass in Fig. 3, Scenario 2. The process is described in the following.
@@ II-E1 Attack on PoI
    At this point, $M_{2}$ would not be able to use PoI, as $C_{1}$ is in $L$, therefore PoI would fail, as shown in Scenario 1
@@
    However, $M_{2}$’s withdrawal is indistinguishable from any other output from the mixer (as long as PoI is not applied) and so cannot be classified as malicious
@@
    So, $M_{2}$ then sends their new funds to the mixer and creates a new deposit with commitment $C_{2}$
@@
    Finally, $M_{3}$ successfully generates a valid PoI, as their withdrawal with commitment $C_{2}$ does not correspond to any of the deposits contained in list $L$
@@
II-E2 Extensions of PoI
There are some recent extensions to PoI. Privacy Pools [36] is a similar protocol to PoI that allows both inclusion and exclusion proofs. Derecho [37] is another similar protocol to PoI that supports proofs for mixers that allow in-pool transactions (also known as “shielded” transactions), i.e., transfers of funds that do not leave the mixer. Derecho uses proof-carrying data for recursive proof composition to achieve this.
However, using Privacy Pools or Derecho does not eliminate the above attack (shown in Fig. 3, Scenario 2). In Privacy Pools, every transaction that originates from the mixer would have to be added to the exclusion list unless a proof is made available to all other participants (e.g., published on chain). Similarly, Derecho has no way to track funds that are withdrawn from a mixer, so the same attack can be implemented.
II-E3 PoI in Practice
@@ II-F Research Aim
We have identified that mixers are commonly used by criminals to launder proceeds of crime (Section II-A); and we have shown that current PoI systems are insufficient, as they can be easily defeated (Section II-E). Therefore, if we want to keep the benefits of anonymity provided by cryptocurrency mixers but also avoid their abuse by bad actors, it is clear that a new design is necessary. To achieve this, we propose a novel design that integrates innocence checks into the mixer (rather than as a separate component, as with current PoI systems); all checks are performed before funds are added to the mixer; and all deposits are verified. This approach ensures that all mixer outputs can be trusted; eliminating the threat of attackers using the mixer to trick the PoI system (i.e., Fig. 3 Scenario 2).
III System Design and Implementation
III-A System Rationale
Existing protocols treat the PoI as a non-compulsory part of the mixer. Therefore, one could attempt to solve the attack vector shown in Fig. 3 Scenario 2 by integrating the PoI system into the mixer. However, this will not completely solve the issue as a malicious user could deposit funds into the mixer, withdraw the funds (with an invalid PoI), deposit the funds again, and finally withdraw the funds to a third address. In such cases, it is not clear whether the second withdrawal will produce an invalid PoI. If the list used by the PoI system is not updated quickly enough, a valid PoI will be produced. Therefore, to ensure a reliable system, the operator must be able to act quickly, which is difficult to achieve in practice.
One could attempt to solve this problem by using instantaneous list updates when invalid PoIs are produced. This is possible as the PoI system is now integrated in the mixer’s smart contract, so the receiver of a withdrawal with an invalid PoI can be added to the exclusion list of the PoI. However, this does not solve the issue, as a malicious user could first transfer the funds to a new address before performing the second deposit. As the mixer smart contract does not have access to transactions that take place outside the contract, this new address cannot be added to the list automatically.
One could attempt to mitigate this by making deposits a two-step process to allow for list updates. In this way, when a deposit arrives, it will not be immediately added to the mixer; rather it is held in a staging area for some predefined time to give the operator of the system time to check if the deposit should be added to the PoI list. Then, after the predefined time has elapsed, the user can add their deposit to the mixer. The rest of the mixer operations should work as before.
This approach solves the attack we have identified in Fig. 3 Scenario 2. However, note that since we have described a system in which all deposits in the mixer can be verified, PoI checks are no longer needed. This is a key idea of our system design, which we introduce in the following sections.
III-B Improvements on ZK Mixer Design
@@
Our protocol improves the design of ZK mixers in three main ways. First, we prevalidate the deposits before they enter the mixer to ensure that they adhere to a set of predefined rules that are configured when the mixer instance is first created. Rules can range from very simple checks that can be performed automatically by smart contracts, such as “only admit deposits from registered users”, to more complex schemes that check the origin of funds or check a user’s historical interactions with other mixers. These latter rules are too complex for a smart contract to perform, and so a group of verifiers is required. The verifiers check every input to the mixer and ensure that it follows the rules of the specific protocol instance. This makes the system highly customizable and eliminates the PoI issue (described earlier in Section II-E).
@@
Second, we introduce a timer delay lock before tokens are added to the Merkle tree of the mixer. The delay is designed to provide verifiers with enough time to verify each input: it can be set as a constant, e.g., one day; or it can be set to vary dynamically, e.g., increasing with system traffic (see Fig. 4 for available delay modes). To be effective, a constant delay must be long enough to ensure that malicious activity can be detected during periods of high traffic; while a variable delay enables both efficient mixer operation during low traffic and sufficient time for validation during high traffic. High traffic bursts over a short period can themselves be an indication of laundering, so dynamically slowing mixer operation during these times acts as a further safety valve. An unusually large number of deposits will trigger a marked increase in the delay time, which flags anomalous deposit behaviour.
@@
Third, we allow verifiers to control inputs to the mixer during the lock period. Using a common consensus protocol, verifiers have the option of admitting deposits, rejecting and returning deposits to the sender, or confiscating deposits. Although confiscation should be used to restitute funds to victims of crime, it can be abused if the verifiers collude. Therefore, when a system instance is first deployed, the confiscation functionality can be enabled or disabled: when disabled, the verifiers only have the option to admit the deposit or reject and return the deposit to the sender. Here, we must stress again that verifiers can only assume control over deposits during the lock period. Once a deposit is accepted into the mixer, it is controlled by the user as in any other ZK-based mixer. Many smart contracts commonly used in DeFi have similar, or even stronger, control mechanisms. For example, the issuers of the most popular stablecoins such as USDC, USDT, and BUSD have the power to freeze or burn stablecoins and do so regularly when ordered by authorities [41]. This property of our system might not satisfy maximalist notions of decentralization, but we believe that it is sufficient for practical use in most cases.
These three additions solve the attack we have identified in Fig. 3 Scenario 2. The attack relies on the ability to pass funds through the mixer to obfuscate their origin, before passing them through the mixer again to circumvent PoI checks. Unlike previous approaches, where the output of the mixer cannot be trusted, in our proposed approach we can trust that withdrawals follow the rules of the mixer. Therefore, it is not possible to perform the attack.
III-C System Components